ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 203, 2014

Argyros, S. A. and P. Motakis A hereditarily indecomposable Banach space with rich spreading model structure

341

Bourgain, J. On the Hardy–Littlewood maximal function for the cube

275

Conlon, D., W. T. Gowers, W. Samotij and M. Schacht On the KŁR conjecture in random graphs

535

Fonf, V. P. and P. Wojtaszczyk Characteristic properties of the Gurariy space

109

Freeman, D., E. Odell, Th. Schlumprecht and A. Zsák Unconditional structures of translates for L_{p}(ℝ^{d})

189

Giannopoulos, A., G. Paouris and B.-H. Vritsiou The isotropic position and the reverse Santaló inequality

1

Gowers, W. T. See D. Conlon Johnson, W. B. and G. Schechtman Subspaces of L_{p} that embed

into L_{p}(μ) with μ finite

211

Johnson, W. B. and A. Szankowski The trace formula in Banach spaces

389

Klartag, B. Concentration of measures supported on the cube  59

König, H. On the best constants in the Khintchine inequality for Steinhaus variables

23

König, H. and V. Milman The derivation equation for C^{k}-functions: stability and localization

405

Lafforgue, V. and A. Naor Vertical versus horizontal Poincaré inequalities on the Heisenberg group

309

Larsen, M. and A. Lindenstrauss Coxeter cochain complexes  173

Latała, R. L_{1}-norm of combinations of products of independent random variables

295

Lindenstrauss, A. See M. Larsen Lindenstrauss, E. and G. Margulis Effective estimates on indefi-

nite ternary forms

445

Litvak, A. E., M. Rudelson and N. Tomczak-Jaegermann On approximation by projections of polytopes with few facets

141

Margulis, G. See E. Lindenstrauss Milman, V. See H. König Motakis, P. See S. A. Argyros Mykhaylyuk, V., M. Popov, B. Randrianantoanina and G.

Schechtman Narrow and ℓ_{2}-strictly singular operators from L_{p}

81

Naor, A. and G. Schechtman On the mathematical contributions of Joram Lindenstrauss

ix

Naor, A. See V. Lafforgue Odell, E. See D. Freeman Oleszkiewicz, K. Precise moment and tail bounds for Rademacher

sums in terms of weak parameters

429

Paouris, G. See A. Giannopoulos

Pisier, G. Random matrices and subexponential operator spaces  223

Popov, M. See V. Mykhaylyuk Preiss, D. Gâteaux differentiability of cone-monotone and pointwise

Lipschitz functions

501

Randrianantoanina, B. See V. Mykhaylyuk Rudelson, M. See A. E. Litvak Samotij, W. See D. Conlon Schacht, M. See D. Conlon Schechtman, G. See W. B. Johnson Schechtman, G. See V. Mykhaylyuk Schechtman, G. See A. Naor Schlumprecht, Th. See D. Freeman Szankowski, A. See W. B. Johnson Tomczak-Jaegermann, N. See A. E. Litvak

Tsirelson, B. Moderate deviations on different scales: no relations  161

Vritsiou, B.-H. See A. Giannopoulos Wojtaszczyk, P. See V. P. Fonf Zsák, A. See D. Freeman